Draw the hyperbola with the equation: 3𝑥 2 − 2𝑦 2 + 6𝑥 − 8𝑦 = 6


1
Expert's answer
2021-02-08T17:33:29-0500

first of all, we need to transform the equation to the canonical form

"\boxed{\frac {(x-x_0)^2} {a^2} - \frac{(y-y_0)^2} {b^2} = 1}" , "x_0" and "y_0" are coordinates of the center,


"3x^{2}+6x-2y^{2}-8y=6"


"(3x^2 + 6x + 3) - (2y^2 +8y+ 8) = 1"


"\dfrac{(x+1)^2}{\dfrac1 3} - \dfrac{(y+2)^2}{\dfrac 1 2} = 1"


O (-1,-2) is the center

"a = \sqrt{\dfrac1 3} \thickapprox 0.6"

"b = \sqrt{\dfrac1 2} \thickapprox 0.7"


what is a and b:



(instead of the origin, there can be any point that is the center of the hyperbola)